OpenStudy (anonymous):

Write the equation of least degree whose roots are -3,-4 and 2

OpenStudy (anonymous):

x=-3 ; x=-4; x=2 x+3 =0; x+4=0; x-2=0 (x+3)(x+4)(x-2) =0, Expand it will become \[x^{3}+5x^{2}-2x-24=0\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

You're welcome !